An object weighing 336 N in air is immersed in water after being tied to a string connected to a balance. The scale now reads 256 N. Immersed in oil, the object appears to weigh 277 N.

(a) Find the density of the object

_______ kg/m3

(b) Find the density of the oil.

________kg/m3
Answer #1

Gravitational acceleration = g = 9.81 m/s2

Density of water = w = 1000 kg/m3

Density of oil =  o

Density of object =  

Weight of object in air = W1 = 336 N

Scale reading when object is immersed in water = W2 = 256 N

Scale reading when object is immersed in oil = W3 = 277 N

Volume of the object = V

When the object is immersed in water,

W2 = W1 - wVg

256 = 336 - (1000)(9.81)V

V = 8.155 x 10-3 m3

W1 = Vg

336 = (8.155x10-3)(9.81)

= 4199.97 kg/m3

When the object is immersed in oil,

W3 = W1 - oVg

277 = 336 - o(8.155x10-3)(9.81)

o = 737.49 kg/m3

a) Density of the object = 4199.97 kg/m3

b) Density of oil = 737.49 kg/m3